# STree: Speculative Tree Decoding for Hybrid State-Space Models

### v1

@@ Abstract
Speculative decoding is a technique to leverage hardware concurrency to improve the efficiency of large-scale autoregressive (AR) Transformer models by enabling multiple steps of token generation in a single forward pass. State-space models (SSMs) are already more efficient than AR Transformers, since their state summarizes all past data with no need to cache or re-process tokens in the sliding window context. However, their state can also comprise thousands of tokens; so, speculative decoding has recently been extended to SSMs. Existing approaches, however, do not leverage the tree-based verification methods, since current SSMs lack the means to compute a token tree efficiently. We propose the first scalable algorithm to perform tree-based speculative decoding in state-space models (SSMs) and hybrid architectures of SSMs and Transformer layers. We exploit the structure of accumulated state transition matrices to facilitate tree-based speculative decoding with minimal overhead to current SSM state update implementations. With the algorithm, we describe a hardware-aware implementation that improves naive application of AR Transformer tree-based speculative decoding methods to SSMs. Furthermore, we outperform vanilla speculative decoding with SSMs even with a baseline drafting model and tree structure on three different benchmarks, opening up opportunities for further speed up with SSM and hybrid model inference. Code will be released upon paper acceptance.
@@ 1 Introduction
Recursive sequence models, such as autoregressive (AR) Transformers, produce a single token with each forward pass. Speculative decoding is a technique to make this process more efficient by leveraging a smaller ‘draft model’ to generate multiple tokens, and a separate ‘verifier model’ to validate the proposed drafts leviathan2022fast . The efficiency stems from exploiting the concurrency of parallel hardware compute to verify multiple tokens in a single model call, while ensuring that the accepted drafts are identical to those that would have been generated by repeated calls to the original model. AR Transformers leverage a sliding window of input tokens (context) as their ‘state’, which is ideal for speculative decoding since their context can be easily edited to remove unverified tokens.
@@
However, modern SSM realizations mamba2 ; mamba are designed to scan through all tokens in the past causally to obtain the state, lacking a mechanism to specify the tree structure. As a consequence, current SSMs can only use the unrolled input (Fig. 1 Left), leading to inefficient repeated token generation. Moreover, SSMs require one state per input sequence in the batch, leading to an explosion of the memory footprint when using the unrolled input. As the size of the tree grows, unrolling the tree quickly becomes infeasible. Therefore, it behooves the designer to look for ways to leverage the characteristics of the hardware to perform multiple steps of state update following the tree structure through a single pass through the model.
Figure 1: Methods to decode a prefix token tree with AR Transformers. A prefix token tree can be unrolled into multiple sequences (used by current SSMs) and computed as a batch or packed into one sequence and use a mask to indicate the tree structure (first used by SpecInfer miao2023specinfer and enabled on SSM by us). The former leads to inefficiency due to repeatedly computed tokens (in orange).
We propose State-Space Speculative Tree Decoding (STree), a method to facilitate tree-based speculative decoding in SSMs through accumulating state transition matrices according to the tree structure. The tree is easily computed at minimal overhead to current SSM state update implementations. We describe a hardware-aware implementation that already in its simplest instantiation shows improvement over the baseline naive application of tree-based speculative decoding to SSMs.
Our contributions in this paper are to (i) propose what, to the best of our knowledge, is the first scalable method to leverage tree decoding in the speculative decoding for both SSMs and hybrid architectures; we also (ii) provide a simplified analysis of the trade-off between acceptance length and model runtime to help determine whether we should scale tree size or even use tree decoding. Finally, we (iii) empirically demonstrate that with a baseline drafting model and static tree structure, there are already improvements in generation speed, thus opening the door to further investigation of possible improvements with more advanced speculative decoding methods employed with transformers.
@@ 2 Related Work
are sequence models that maintain a hidden state and use it to predict the next datum in the sequence given the previous observations in the sequence lindquist1979stochastic . Such models can be stacked, so the output of one is the input of another s4 ; zancato2022stacked , and their parameters can be input-dependent krener1975bilinear ; mamba . In order to scale model size, transition matrices are often restricted to be diagonal mamba2 ; de2024griffin , while to leverage efficient hardware-aware implementations, SSM layers are often interleaved with Transformers in Hybrid architectures lieber2024jamba ; glorioso2024zamba , also useful for initial distillation wang2024mamba . Our method is based on scalable SSMs of the form (1) for cleanliness of presentation. In the actual implementation, $A$ and $B$ are discretized by $\Delta$ and the $y_{t}$ is also additionally dependent on $D(u_{t})u_{t}$.
$\begin{cases}x_{t+1}=A(u_{t})x_{t}+B(u_{t})u_{t}\\ y_{t}=C(u_{t})x_{t}\end{cases}$ (1)
@@
denotes a set of techniques designed to parallelize sequential inference in large language models (LLMs) by utilizing a smaller ‘draft model’ to produce multiple candidate trajectories and a large ‘verifier model’ to test them for consistency with the original model (leviathan2022fast, ). Sub-networks can also be used for drafting in a form of ‘self-speculation’ (liu2024kangaroo, ), while speculation can be applied to latent features (li2024eagle, ; li2024eagle2, ) in addition to the output. Structural changes to the speculative processes include tree-based verification (miao2024specinfer, ; chen2024sequoia, ), multi-head decoding (cai2024medusa, ), and beam sampling(qin2025beam, ; qin2025mtjd, ). The same methods can also be used for ‘lossy verification’ using judge decoding (bachmann2025judge, ).
@@
miao2023specinfer has been shown effective in improving the acceptance length of the drafted sequence li2024eagle ; li2024eagle2 ; chen2024sequoia ; cai2024medusa ; svirschevski2024specexec ; qin2025beam ; qin2025mtjd . EAGLE li2024eagle reported a increase in acceptance length by using a static tree structure, leading to an overall speed up ratio improvement. Sequoia chen2024sequoia generates the optimal tree for the speculated tokens using a dynamic programming algorithm and achieved further improvement. The underlying mechanism that enables tree-based verification is transformer’s ability to specify which tokens to attend to with an arbitrary attention mask individually for each token. A prefix token tree can be packed into a sequence and a topology aware mask miao2023specinfer can be used to inform the self attention block the tree structure.
@@
To the best of our knowledge, our work is the first to propose a scalable algorithm to leverage tree-based verification to perform speculative decoding in SSMs as well as hybrid architectures. Using a baseline drafting model and an elementary tree structure, we show that SSMs and hybrid models can benefit from speculative tree decoding.
@@ 3 Method
Given a prefix token tree $T$ with tokens $\{t_{1},\dots,t_{N}\}$ as vertices and $t_{1}$ as the root node, we can pack the tokens into one sequence $S=\{t_{1},\dots,t_{N}\}$ and represent the topology of the tree using a special attention mask $L\in\{0,1\}^{N\times N}$ (‘tree’ mask). Each vertex in $T$ has one unique path to the root node, and we denote these paths as $s_{i}=\{t_{n}|t_{n}\text{ is in the path from }t_{1}\text{ to }t_{i}\}$. Then the tree mask $L$ can be constructed by the indicator function:
$L_{i,j}=\mathbbm{1}_{s_{i}}\{t_{j}\}$ (2)
@@
The input features $u$ is first projected into the parameters using a linear projection. Here we abuse the notation to let $X_{t}:=X(u_{t})$:
@@
Following the simplified SSM in Eqn.(1), for each $y_{t}=C_{t}x_{t}$, we can expand it recursively into:
@@
Same as mamba ; mamba2 ; de2024griffin ; lieber2024jamba , we enforce a diagonal structure on $A_{i}$ to reduce the products in Eqn.(4) to a sum of logarithms:
@@
Since $A_{i}$ is a diagonal matrix, we can extract the diagonal of $\log A_{i}$ as a vector and put them into a new matrix $A_{log}=[diag(\log A_{1}),\dots,diag(\log A_{N})]^{T}\in\mathbb{R}^{N\times d_{x}}$. We define:
@@
Based on the methods presented in previous paragraphs, we propose an algorithm in Alg. 1 to perform State-space speculative decoding with tree decoding. The algorithm centers on a tree scan kernel that computes the output of the output of a packed prefix tree input and also caches the intermediate activation values (i.e. $A_{i:j},B_{i:j},C_{i:j}$):
@@
[…]
@@
5:     while $should\_continue$ do
6:         $L_{i:j},t_{i:j}\leftarrow\textsc{Draft}(t_{i-1})$triangleright Draft a tree with last accepted token
7:         $x^{*}\leftarrow\textsc{ActivationReplay}(L^{*},Cache)$triangleright Recompute state up to the rejected tokens
8:         $t^{\prime}_{i:j},Cache\leftarrow\textsc{TreeScan}(L,t_{i:j},x^{*})$triangleright Getting output and cache from target model
9:         $L^{*},t_{i:k}\leftarrow\textsc{FirstRejected}(t^{\prime}_{i:j},t_{i:j})$triangleright Accept/Reject Drafted tokens
@@ 4 Analysis
The success of speculative decoding relies on the fact that computing an input of length $N>1$ doesn’t increase $t$ too much. As SSMs are not as good as transformers in performing parallel computation due to their recurrent nature, the runtime of the target model calls sees a non-negligible increase even at small input length. Therefore, there exists a trade-off between the gain in acceptance length and the increase in runtime of target model calls as the size of the tree grows. Given that SSM compute and memory cost is linear in the input length, we assume a simple linear model between the model runtime and the input size, where $t=kN+C$, where $C$ is a constant contributed by the time for loading model weights and $kN$ is proportional to input size $N$, contributed by loading inputs and computation. When the models are large, constant $C$ dominates the runtime, leading to a negligible effect on runtime by input length $N$ when $N$ is small.
Figure 2: Left: The runtime for a call to Mamba2-2.7B model vs. input size with STree and Fuse Selective Scan (FSS). A linear regression is performed to obtain the gradient and interception. Middle: The runtime for a call to MambaInLlama-8B model vs. input size with STree and FSS. A polynomial regression with degree of 2 is used to obtain the parameters. Right: The ratio of acceptance length $\tau$ required to get wall time improvement vs. input length $N_{2}$, with a fixed $N_{1}=5$
Therefore, to gain wall time improvement we require:
@@
To determine whether we should use tree scan for speculative decoding, we want to evaluate whether the gain in acceptance length can outweigh the increase in runtime. For example, we measured the runtime for Mamba2-2.7B and MambaInLlama-8B model calls using both STree and Fused Selective Scan (FSS), which is an algorithm optimized for short sequence inference wu2024snakes , with different input sizes in Fig. 2(a) and Fig. 2(b). We performed linear and quadratic regression to obtain an approximation of the parameters ($k,C,a,b,c$). Assuming that we are comparing STree against vanilla state-space speculative decoding that uses FSS to compute forward pass where we draft 1 sequences with 5 tokens, we can let $N_{1}=5$ and plot the ratio of acceptance length $\tau_{2}/\tau_{1}$ required to achieve wall time improvement for both models in Fig. 2(c). We can see that for STree with Mamba2-2.7B, to gain a wall time improvement at a tree size $N_{2}$ of 15 tokens, we need the drafted tree to achieve at least 1.5×1.5$\times$1.5 × acceptance rate as compared to vanilla speculative decoding. On the other hand, STree with Mamba2InLlama-8B requires at least $\times1$ acceptance rate with the same tree size, which is much easier to achieve than Mamba2-2.7B. This means that the performance gain that can be achieved with tree decoding with Mamba2InLlama-8B is much more than that with Mamba2-2.7B. This could be due to Mamba2InLlama-8B being a larger model with a larger constant overhead for model call.
@@ 5 Experimental Results
In this section, we aim to demonstrate the efficiency of STree on speculative decoding. In Sec. 5.1, we compare STree against the unrolled input baseline to show that STree is more efficient at decoding a tree with SSM. Then, in Sec. 5.2, we compare against speculative decoding without tree-based verification to show that STree is able to achieve speed improvement with baseline drafting models and tree construction methods. All experiments are run on an Nvidia RTX 3090 GPU.
@@ 5.1 Efficiency of STree against unrolled baseline
STree performs on par with FSS at small tree size (4-layer), as FSS is optimized for short sequences. However, as the size of the tree increases, the forward pass with STree increases slightly from $27.6$ms to $34.0$ms, while FSS increases drastically from $27.3$ms to $59.8$ms. This is due to unrolling a prefix tree introduced repeated tokens and extra states as shown in Fig. 3 Left. This slows down the forward pass for both FSS as well as chunk scan. As STree is able to directly decode a packed tree sequence, it demonstrates good scalability as the input tree size grows because it avoids repeated computations. Meanwhile, forward pass with chunk scan is consistently slower than both fused selective scan and STree, as it is optimized for parallel training for long sequences and pays a big overhead at a short sequence length wu2024snakes ; wang2024mamba . Hence, it should not be considered for the use of speculative decoding.
Tree depth Methods #. of States Tokens Computed

### v2

@@ Abstract
Speculative decoding is a technique to leverage hardware concurrency in order to enable multiple steps of token generation in a single forward pass, thus improving the efficiency of large-scale autoregressive (AR) Transformer models. State-space models (SSMs) are already more efficient than AR Transformers, since their state summarizes all past data with no need to cache or re-process tokens in the sliding window context. However, their state can also comprise thousands of tokens; so, speculative decoding has recently been extended to SSMs. Existing approaches, however, do not leverage the tree-based verification methods, since current SSMs lack the means to compute a token tree efficiently. We propose the first scalable algorithm to perform tree-based speculative decoding in state-space models (SSMs) and hybrid architectures of SSMs and Transformer layers. We exploit the structure of accumulated state transition matrices to facilitate tree-based speculative decoding with minimal overhead relative to current SSM implementations. Along with the algorithm, we describe a hardware-aware implementation that improves naive application of AR Transformer tree-based speculative decoding methods to SSMs. Furthermore, we outperform vanilla speculative decoding with SSMs even with a baseline drafting model and tree structure on three different benchmarks, opening up opportunities for further speed up with SSM and hybrid model inference. Code can be find at: https://github.com/wyc1997/stree.
@@ 1 Introduction
Recursive sequence models, such as autoregressive (AR) Transformers, produce a single token with each forward pass. Speculative decoding is a technique to make this process more efficient by leveraging a smaller ‘draft model’ to generate multiple tokens, and a separate ‘verifier model’ to validate the proposed drafts leviathan2023Fast . The efficiency stems from exploiting the concurrency of parallel computer hardware to verify multiple tokens in a single model call, while ensuring that the accepted drafts are identical to those that would have been generated by repeated calls to the original model. AR Transformers leverage a sliding window of input tokens (context) as their ‘state’, which is ideal for speculative decoding since their context can be easily edited to remove unverified tokens.
@@
However, modern SSM realizations mamba2 ; mamba are designed to scan through all tokens in the past causally to obtain the state, lacking a mechanism to specify the tree structure. As a consequence, current SSMs can only use the unrolled input (Fig. 1 Left), leading to inefficient repeated token generation. Moreover, SSMs require one state per input sequence in the batch, leading to an explosion of the memory footprint when using the unrolled input. As the size of the tree grows, unrolling the tree quickly becomes infeasible. Therefore, we seek ways to leverage the characteristics of the hardware to perform multiple steps of state update following the tree structure through a single pass through the model.
Figure 1: Methods to decode a prefix token tree with AR Transformers. A prefix token tree can be unrolled into multiple sequences (used by current SSMs) and computed as a batch or packed into one sequence using a mask to indicate the tree structure (first used by SpecInfer miao2023specinfer and extended to SSMs here). The former leads to inefficiency due to repeatedly computed tokens (in orange).
We propose State-Space Speculative Tree (STree) decoding, a method to facilitate tree-based speculative decoding in SSMs through accumulating state transition matrices according to the tree structure. The tree is easily computed with minimal overhead relative to current SSM state update implementations. We describe a hardware-aware implementation that already in its simplest instantiation shows improvement over the baseline naive application of tree-based speculative decoding to SSMs.
Our contributions in this paper are to (i) propose what, to the best of our knowledge, is the first scalable method to leverage tree decoding in the speculative decoding for both SSMs and hybrid architectures; we also (ii) provide a simplified analysis of the trade-off between acceptance length and model runtime to help determine whether we should scale tree size or even use tree decoding. Finally, we (iii) empirically demonstrate that with a baseline drafting model and static tree structure, there are already improvements in generation speed, thus opening the door to further investigation of more advanced speculative decoding methods employed with transformers.
@@ 2 Related Work
are sequence models that maintain a hidden state and use it to predict the next datum in the sequence given the previous ones lindquist1979stochastic . Such models can be stacked, so the output of one is the input of another s4 ; zancato2022stacked , and their parameters can be input-dependent krener1975bilinear ; mamba ; zancato2024b . In order to scale model size, transition matrices are often restricted to be diagonal mamba2 ; de2024griffin , while to leverage efficient hardware-aware implementations, SSM layers are often interleaved with Transformers in Hybrid architectures lieber2024jamba ; glorioso2024zamba , also useful for initial distillation wang2024mamba . Our method is based on scalable SSMs of the general form (1).
$\begin{cases}x_{t+1}=A(u_{t})x_{t}+B(u_{t})u_{t}\\ y_{t}=C(u_{t})x_{t}+D(u_{t})u_{t}\end{cases}$ (1)
@@
denotes a set of techniques designed to parallelize sequential inference in large language models (LLMs) by utilizing a smaller ‘draft model’ to produce multiple candidate trajectories and a large ‘verifier model’ to test them for consistency with the original model (leviathan2023Fast, ). Sub-networks can also be used for drafting, a form of ‘self-speculation’ (liu2024kangaroo, ), while speculation can be applied to latent features (li2024eagle, ; li2024eagle2, ) in addition to the output. Structural changes to the speculative processes include tree-based verification (miao2024specinfer, ; chen2024sequoia, ), multi-head decoding (cai2024medusa, ), and beam sampling(qin2025beam, ; qin2025mtjd, ). The same methods can also be used for ‘lossy verification’ using judge decoding (bachmann2025judge, ).
@@
miao2023specinfer has been shown effective in improving the acceptance length of the drafted sequence li2024eagle ; li2024eagle2 ; chen2024sequoia ; cai2024medusa ; svirschevski2024specexec ; qin2025beam ; qin2025mtjd . EAGLE li2024eagle reported an increase in acceptance length by using a static tree structure, leading to an overall speed up ratio improvement. Sequoia chen2024sequoia generates the optimal tree for the speculated tokens using a dynamic programming algorithm and achieved further improvement. The underlying mechanism that enables tree-based verification is Transformers’ ability to specify which tokens to attend to with an arbitrary attention mask individually for each token. A prefix token tree can be packed into a sequence and a topology aware mask miao2023specinfer can be used to inform the self attention block the tree structure.
@@ 3 Method
[…]
$L_{i,j}=\mathbbm{1}_{s_{i}}\{t_{j}\}.$ (2)
@@
The input feature $u$ is first mapped onto the parameters using a linear projection. Here we abuse the notation to let $X_{t}:=X(u_{t})$:
@@
Ignoring the last term $D_{t}$ in Eqn.(1) for simplicity of notation, for each $y_{t}=C_{t}x_{t}$, we can expand it recursively into:
@@
As done in mamba ; mamba2 ; de2024griffin ; lieber2024jamba , we enforce a diagonal structure on $A_{i}$ to reduce the products in Eqn.(4) to a sum of logarithms:
@@
Since $A_{i}$ is diagonal, we can represent the diagonal of $\log A_{i}$ as a vector and assemble a new matrix $A_{log}=[diag(\log A_{1}),\dots,diag(\log A_{N})]^{T}\in\mathbb{R}^{N\times d_{x}}$. We define:
@@
Based on the methods presented in previous paragraphs, we propose an algorithm in Alg. 1 to perform State-space speculative decoding with tree decoding. The algorithm centers on a tree scan kernel that computes the output of a packed prefix tree input and also caches the intermediate activation values (i.e., $A_{i:j},B_{i:j},C_{i:j}$):
@@
3:  Initialize $Cache:\text{activation cache to facilitate recomputation of state}$
@@
5:  while should_continueshould\_continue do
6:   $L_{i:j},t_{i:j}\leftarrow\textsc{Draft}(t_{i-1})$ $\triangleright$ Draft a tree with last accepted token
7:   $x^{*}\leftarrow\textsc{ActivationReplay}(L^{*},Cache)$ $\triangleright$ Recompute state up to the rejected tokens
8:   $t^{\prime}_{i:j},Cache\leftarrow\textsc{TreeScan}(L,t_{i:j},x^{*})$ $\triangleright$ Getting output and cache from target model
9:   $L^{*},t_{i:k}\leftarrow\textsc{FirstRejected}(t^{\prime}_{i:j},t_{i:j})$ $\triangleright$ Accept/Reject Drafted tokens
@@ 4 Analysis
The success of speculative decoding relies on the fact that computing an input of length $N>1$ doesn’t increase $t$ too much. As SSMs are not as good as transformers in performing parallel computation due to their recurrent nature, the runtime of the target model calls sees a non-negligible increase even at small input length. Therefore, there exists a trade-off between the gain in acceptance length and the increase in runtime of target model calls as the size of the tree grows. Given that SSM compute and memory cost is linear in the input length, we assume a simple linear model between the model runtime and the input size, where $t=kN+C$, where $C$ is a constant contributed by the time for loading model weights and $kN$ is proportional to input size $N$, contributed by loading inputs and computation. When the models are large, the constant $C$ dominates the runtime, leading to a negligible effect on runtime by input length $N$ when $N$ is small.
Figure 2: Left: The runtime for a call to Mamba2-2.7B model vs. input size with STree and Fuse Selective Scan (FSS). A linear regression is performed to obtain the slope and intercept. Middle: The runtime for a call to MambaInLlama-8B model vs. input size with STree and FSS. A polynomial regression with degree 2 is used to obtain the parameters. Right: The ratio of acceptance length $\tau$ required to get wall-clock time improvement vs. input length $N_{2}$, with a fixed $N_{1}=5$
Therefore, to gain wall-clock time improvement we require:
@@
To determine whether we should use tree scan for speculative decoding, we want to evaluate whether the gain in acceptance length can outweigh the increase in runtime. For example, we measured the runtime for Mamba2-2.7B and MambaInLlama-8B model calls using both STree and Fused Selective Scan (FSS), which is an algorithm optimized for short sequence inference wu2024snakes , with different input sizes in Fig. 2(a) and Fig. 2(b). We performed linear and quadratic regression to obtain an approximation of the parameters ($k,C,a,b,c$). Assuming that we are comparing STree against vanilla state-space speculative decoding that uses FSS to compute forward pass where we draft 1 sequence with 5 tokens, we can let $N_{1}=5$ and plot the ratio of acceptance length $\tau_{2}/\tau_{1}$ required to achieve wall-clock time improvement for both models in Fig. 2(c). We can see that for STree with Mamba2-2.7B, to gain a wall-clock time improvement at a tree size $N_{2}$ of 15 tokens, we need the drafted tree to achieve at least $1.5\times$ acceptance rate as compared to vanilla speculative decoding. On the other hand, STree with Mamba2InLlama-8B requires at least $1.1\times$ acceptance rate with the same tree size, which is much easier to achieve than Mamba2-2.7B. This means that the performance gain that can be achieved with tree decoding with Mamba2InLlama-8B is much more than that with Mamba2-2.7B. This could be due to Mamba2InLlama-8B being a larger model with a larger constant overhead for model call.
Models of bigger size
Theoretically, the effectiveness of speculative decoding depends on the runtime difference between the draft model and the target model as well as the acceptance rate of the drafted tokens. Everything else being equal, using a bigger target model leads to a bigger runtime difference. Therefore, the effectiveness of our method on bigger models should hold or improve. In Table 1, we provide the results of latency of a Mamba2 forward pass using STree vs. autoregressive forward with selective scan with same input tree configuration. Note that for 7B, 13B, and 23B models, we initialize the model with random weight to measure the runtime since there is no pretrained model of those sizes. We can see that as the model size grows, the relative overhead of using STree decreases from 2.04x to 1.26x, signaling that as model sizes increase, our method is likely going to be even faster, given that the average acceptance length stays the same. Meanwhile, the gap between the Vanilla SD and STree is closing as model size increases (from 48.9% with 2.3B to 18.9% with 23B), which is another signal that our method is scalable to larger models.
Table 1: Latency of a Mamba2 forward pass using STree vs. autoregressive forward with selective scan. Note that for 7B, 13B, and 23B models, we initialize the model with random weight to measure the runtime since there is no pretrained model of those sizes.
2.3B 7B 13B 23B
Autoregressive 10.95 22.94 40.69 72.41
STree 22.36 (2.04x) 33.79 (1.47x) 55.90 (1.37x) 91.08 (1.26x)
Vanilla SD 15.05 (1.37x) 26.46 (1.15x) 45.95 (1.13x) 76.74 (1.06x)
@@ 5 Experimental Results
In this section, we aim to demonstrate the efficiency of STree on speculative decoding. In Sec. 5.1, we compare STree against the unrolled input baseline to show that STree is more efficient at decoding a tree with SSMs. Then, in Sec. 5.2, we compare against speculative decoding without tree-based verification to show that STree is able to achieve speed improvement with baseline drafting models and tree construction methods. All experiments are run on an Nvidia RTX 3090 GPU.
@@ 5.1 Efficiency of STree against unrolled baseline
STree performs on par with FSS at small tree size (4-layer), as FSS is optimized for short sequences. However, as the size of the tree increases, the forward pass with STree increases slightly from 27.627.6ms to 34.034.0ms, while FSS increases drastically from 27.327.3ms to 59.859.8ms. This is due to unrolling a prefix tree introduced repeated tokens and extra states as shown in Fig. 3 Left. This slows down the forward pass for both FSS as well as chunk scan. As STree is able to directly decode a packed tree sequence, it demonstrates good scalability as the input tree size grows because it avoids repeated computations. Meanwhile, forward pass with chunk scan is consistently slower than both fused selective scan and STree, as it is optimized for parallel training for long sequences and pays a big overhead at a short sequence length wu2024snakes ; wang2024mamba . Hence, it should not be considered for use in speculative decoding.
Tree
depth
Methods
#. of
States
Tokens
Computed